A body cools from 70 -60 in 10 minutes 60-56 in 10 minutes. What is the temperature of body in next 10 minutes?

Grade:12

1 Answers

jitender
6 years ago
This ques comes from Physics thermodynamics. Applying Newton law of coolingRate of cooling = ◇T/◇t = k (T-T°)◇= changeT= temp average t= time T°= surrounding Temp 70-60 / 10 = K ( 68- T° ) : T=70+60/2=65K=1/68-T° 1st60-56 /10 = K (58-T°) K=2/5(58-T°) 2ndDivide 1st with 2ndT°= 51.33Now 56-Tt / 10 = 2/5(6.67) [(56+Tt)/2 -51.33]Tt=53.85